# ObjectVLA: End-to-End Open-World Object Manipulation Without Demonstration

### v1

@@ Abstract
Figure 1: Cost-effective object generalization without demonstration: Users capture object photos via smartphones and generate visual-text data to continually learn on a pre-trained VLA model with limited resources.
@@ 1 Introduction
Our approach begins by curating a dataset of visual-textual pairs augmented with localization metadata (e.g., bounding boxes). This dataset is co-finetuned with teleoperated robot interaction data, while the robot data itself is enriched with localization-guided reasoning. By embedding localization as a bridging representation, we create a unified pathway between visual-language inputs and robotic actions. This enables zero-shot object generalization: the model can recognize and manipulate novel objects—even those absent from robot training data—without task-specific retraining.
@@
The versatility of our approach is further demonstrated across diverse scenarios, including bin-picking and tasks requiring composite skills like pushing and rotating. Notably, our framework supports rapid adaptation to novel objects: by collecting smartphone-captured images and performing lightweight fine-tuning, the model generalizes to objects absent from the original dataset. An example is illustrated in Figure 1. These experiments underscore our method’s ability to reduce reliance on large-scale human demonstrations while achieving robust object generalization.
Our primary contribution is a unified pipeline for integrating vision-language datasets with robot interaction data, enabling end-to-end object generalization. Through systematic evaluation, we validate the framework’s performance on complex multi-stage tasks (e.g., bin-picking) and multi-skill manipulation (e.g., rotating, pushing), highlighting its universality. Despite some of the existing works, such as RT-2 [4] and ECoT [24] giving a glimpse of how co-finetuning can achieve simple object generalization, they neither elucidate the underlying mechanism of achieving such generalization nor address the boundary of their methodologies. In contrast, our approach — though simple and straightforward — demonstrates that training VLA models with a hybrid dataset of robot interaction data and visual-textual data significantly enhances generalization. This level of generalization goes significantly beyond previously demonstrated end-to-end approaches. Crucially, our framework enables practical deployment: even a small set of smartphone images and brief fine-tuning suffices to adapt the model to novel objects, significantly advancing real-world robotic flexibility.
@@ 2 Related Work
Vision-language-action models for robot control. Recent research has focused on developing generalist robot policies trained on increasingly expansive robot learning datasets [16, 10]. Vision-language-action models (VLAs) represent a promising approach for training such generalist policies [11, 2, 24]. VLAs adapt vision-language models (VLMs), pre-trained on vast internet-scale image and text data, for robotic control [9, 15]. This approach offers several advantages: leveraging large vision-language model backbones, with billions of parameters, provides the necessary capacity for fitting extensive robot datasets. Furthermore, reusing weights pre-trained on internet-scale data enhances the ability of VLAs to interpret diverse language commands and generalize to novel objects and environments. However, current VLA models struggle to recognize open-world objects when these objects absent from the robot interaction data [11, 23]. This is mainly due to VLMs essentially “overwrites” its previously acquired knowledge of open-world objects with robot-specific information.
Generalization in robot learning. In the realm of robot learning, generalization, particularly object generalization, remains a core challenge and active area of research. Many works leverage techniques such as domain randomization, meta-learning, and data augmentation to improve a robot’s ability to recognize and interact with novel objects unseen during training. For instance, domain randomization methods [20, 8] randomize visual and physical parameters during simulation training to force the agent to learn features invariant to these irrelevant details, leading to better real-world generalization. Furthermore, meta-learning approaches [6] aim to train models that can rapidly adapt to new objects with limited data, directly addressing the object generalization problem. Finally, data augmentation methods [13, 12], enhance the diversity of the training data, exposing the model to a wider range of object appearances and orientations, thereby promoting robustness and generalization to novel objects. There is also a field of work using large language models or vision-language models to do open-vocabulary manipulation [19, 25, 14, 1], combined with motion planning and robot learning methods. However, these approaches involve separate modules that are trained independently for different components. As the best of our knowledge, we are the first work to discuss object generalization beyond certain category in visuomotor policy learning.
@@ 3.1 Notation and Motivation
Given a set of expert demonstrations that contain complex robot skill trajectories, we want to learn a visuomotor policy $\pi:\{\mathcal{O}_{r},\mathcal{I}_{r}\}\mapsto\mathcal{A}$ that maps the visual observations $o_{r}\in\mathcal{O}_{r}$ and the language instruction $i_{r}\in\mathcal{I_{r}}$ to actions $a\in\mathcal{A}$. The action changes accordingly when the language instruction and visual input change. The $r$ denote for data in human demonstration data. Typically, for each language instruction it contains robot skill such as ”push” or ”pick up” and the target object, which is denoted as $\{obj_{r},skill_{r}\}\in i_{r}$. We then formally define the visual-text data, where $\varphi:\{\mathcal{O}_{v},\mathcal{I}_{v}\}\mapsto\mathcal{L}_{v}$, where we input the image $o_{v}\in\mathcal{O}_{v}$ and give a language instruction $i_{v}\in\mathcal{I}_{v}$, the model is output with the corresponding answer $l_{r}\in\mathcal{L}_{v}$. The notation $v$ denotes visual-text data.
In this work, we explore the generalization of objects, focusing on those that are not part of the robot interaction data but are present in visual-textual data.
@@ 3.2 Data Construction
Visual-textual data construction. To explore the model’s ability to generalize to novel object, we constructed a diverse visual-textual dataset. For the visual component, we collected 100 distinct objects that are not included in the robot interaction objects. Specifically, using three cameras mounted on the robot (see Figure 2), we captured 20 images per object, covering various poses and orientations to ensure diversity. For the textual component, we employ a fixed template, “Detecting the bounding box of object.”, as the question, and the corresponding bounding box as the answer. In total, our vision-language dataset comprises 2,000 visual-textual pairs.
Reasoning data construction. We utilize localization metadata to bridge the gap between visual-text data and robot data, as previously mentioned. To establish this implicit link between visual-text and action, we incorporate localization metadata into the robot data. This section details how we construct reasoning with localization for robot data.
For each task, we first identify target objects based on the language instructions. We then employ DinoX [18], a cutting-edge open-vocabulary object detector, to annotate the bounding boxes of these objects. DinoX can generate a bounding box given an object’s name. To ensure accuracy, we manually verify and correct any erroneous bounding boxes produced by DinoX. Since our workspace has two external camera views, which can result in different bounding boxes for the same object, we annotate only one (right camera in our experiments). Following Qwen2-VL [21], we use a fixed template, “¡—object_ref_start—¿object¡—object_ref_end—¿¡—box_start— $(x_{1},y_{1})$,$(x_{2},y_{2})$¡—box_end—¿.”, to represent the localization reasoning. This reasoning is generated before each action and injected into the policy model through a learnable module. For a detailed explanation of this injection module’s architecture, we refer readers to DiVLA [22], the base model used in our experiments. An example of contructed visual-text data is at Figure 3.
@@ 3.3 Training Strategy and Implementation Details
For our experiments, we adopt diffusion-based VLA, a widely used class of Vision-Language-Action (VLA) models exemplified by methods like $\pi_{0}$ [2] and TinyVLA [23]. We select diffusion-based VLA over auto-regressive alternatives due to its significantly faster inference speed, a critical advantage for real-time robotic applications (see FAST [17] for a detailed comparison). Specifically, we utilize DiVLA [22], a representative VLA architecture, co-train on a hybrid dataset comprising robot interaction data and the vision-language corpus. To balance task-specific adaptation and semantic generalization, we maintained a 10:1 data ratio (robot-to-visual-textual data) across all tasks. This ratio empirically proved sufficient for robust object generalization, aligning with prior findings on the benefits of co-training for VLA capabilities. Notably, increasing the proportion of robot data beyond this ratio led to a decline in in-domain task success rates. We hypothesize this stems from the limited capacity of the 2B-parameter DiVLA model compared to larger architectures like ECoT (7B) [24] and RT-2 (55B) [4], which can better absorb domain-specific data without overfitting.
@@ 4 Experiments
Figure 3: Example of constructed visual-text data. Left: Photo taken by the robot’s camera. Right: Object captured with a smartphone.
@@ 4.1.1 Experimental Setup
Ablation study. To further understand our method’s effectiveness, we conducted an ablation study. We found that object generalization relies heavily on two key factors: first, explicitly linking vision and language to action through bounding boxes. This provides a direct connection between the visual object, its linguistic description, and the required manipulation. Second, designing a reasoning process in the robot data that mirrors the structure of vision-language pair data. This allows the model to leverage the rich information encoded in pre-trained vision-language models.
@@ 4.1.2 More Observations
Figure 4: Example Objects Used in Experiments. Left: Objects present in the robot training data. Right: Examples of novel objects, not present in the robot data, but included in the visual-text co-training dataset (see Appendix for a comprehensive list).
@@ 4.2 Combining with More Skills
Implementation details. We train one model for each skill to ensure that the model focuses more on understanding the objects rather than multi-task learning. We use the same visual-textual data of “move” task. Following established protocols from our prior work, this visual-text dataset trained concurrently with the demonstration data for comprehensive evaluation. Each object was tested with 3 trials. In total, 150 trials were conducted. The training setting is provided in Appendix.
Results. As shown in Table 1, our method achieved high success rates on the robot interaction objects for both rotate and push skill. Analysis of the failed “rotate” trials revealed that the primary cause is the model’s inability to grasp the target object securely. When evaluating performance on out-of-distribution (OOD) objects, we observed a decrease in task completion rates compared to in-distribution objects, as expected. However, the model still successfully completed nearly two-thirds of the trials. Notably, in most failure cases, the model did not incorrectly identify the target object but rather failed to execute the skill completely. This was particularly evident in the “rotate” trials, where successful execution hinges on a secure grasp, a challenging requirement for unseen objects. Nevertheless, these experiments strongly support the claim that ObjectVLA can transfer learned skills, beyond basic pick and place, to novel objects within the framework we have developed. The results underscore the potential of ObjectVLA for generalized robotic manipulation, capable of adapting to new objects and tasks beyond its initial training.
@@ 4.3 Instruction-Driven Bin Picking
To further evaluate ObjectVLA, we conducted experiments in a more practical scenario: end-to-end instruction-driven bin-picking. Unlike prior works (e.g., GR-2 [5] and DiVLA [22]) that execute bin-picking tasks without specific semantic instructions—typically limited to generic actions like transferring all objects from one container to another—we focus on a significantly more challenging setting [5, 22]. In our experiments, the robot is required to identify and retrieve a specific target object based on natural language instructions (e.g., ”Pick the hexagonal bolt from the bin”). This scenario elevates the complexity of conventional bin-picking tasks by integrating cross-modal understanding (vision-to-language alignment) and fine-grained object discrimination that have multiple objects in the scene. Notably, the objects are randomly placed on the panel, which is a large area. Not only the model need to figure out the object’s position, but also need to aware of its pose.
Implementation details. We collected new data within this environment. For robot interaction data, we collected 600 pick-and-place trajectories using the same ”seen” objects as in previous experiments. For visual-text data, we used half the number of objects from previous experiments, capturing 20 images of each. We compared our method against OpenVLA, a state-of-the-art VLA model, reporting success rates for both in-distribution and out-of-distribution objects. Evaluation consisted of three trials per object, totaling 183 trials per method. In each trial, at least two objects were randomly placed on the plate, and the model was instructed to pick and place a specific object according to the given instruction.
Results. Table 2 presents our experimental results. Bin picking, requiring object retrieval from random positions and poses, poses a significant challenge even for in-distribution objects. OpenVLA achieves a success rate of only 42.4% for in-distribution objects, significantly less than half. Surprisingly, it still completed roughly 10% of trials with out-of-distribution objects. This is likely due to some test objects sharing attributes with training objects (e.g., bread resembling a muffin, a green mug differing only in color from a brown training mug). In contrast, our method successfully completed 87 of 150 trials, including many completely novel objects, a 46.7% improvement over OpenVLA. This further emphasizes the necessity of co-training with both robot interaction and visual-text data for effective object generalization.
Figure 6: Experimental results for smart-phone captured objects and trained by continual learning. We test two new objects. We take pictures of these two objects via smart phone, and continually trained the objects on pre-trained model. Each object evaluated across 10 trials. We report the success rate for each object.
@@ 4.4 Cheap Object Generalization via Smart-Phone Pictures and Continual Learning
The previous section demonstrated that our proposed co-training strategy enables the imitation learning method to generalize to any object by constructing corresponding visual-text data for each object. This approach significantly enhances the model’s ability to handle a broader range of objects, without requiring extensive retraining. However, there are two key limitations that need to be addressed.
@@
Therefore, in this section, we test a simpler and more cost-effective approach: using a smart-phone camera to collect images from various perspectives. The model is then continuously trained on the pre-trained weights using this more accessible data collection method. As shown in Figure 6, we test with two objects: Pikachu and a brown toy cat. For each object, we capture 21 images and follow the same data construction pipeline discussed earlier. We train these objects in bin-picking environment. Our results demonstrate that the model is able to recognize and successfully grasp the objects with a high success rate, 80% for Pikachu and 90% success rate for toy cat. More importantly, we only need to continue training the model for 1 epoch. Because the collected data size is small, the training process can be extremely fast that can be finished up to ten minutes. This validates the effectiveness of our approach, showing that simple smartphone image collection combined with continuous learning enables open-world object manipulation in an end-to-end model. This experiment demonstrates that our method is flexible and cost-effective, making it a plug-and-play solution for existing VLA models, enabling them to generalize to virtually any object.

### v2

@@ 1 Introduction
Our approach begins by curating a dataset of image-text pairs augmented with localization metadata (e.g., bounding boxes). This dataset is co-finetuned with teleoperated robot interaction data, while the robot data itself is enriched with localization-guided reasoning. By embedding localization as a bridging representation, we create a unified pathway between visual-language inputs and robotic actions. This enables zero-shot object generalization: the model can recognize and manipulate novel objects—even those absent from robot training data—without task-specific retraining.
@@
The versatility of our approach is further demonstrated across diverse scenarios, including bin-picking and tasks requiring composite skills like pushing and rotating. Notably, our framework supports rapid adaptation to novel objects: by collecting smartphone-captured images and performing lightweight fine-tuning, the model generalizes to objects absent from the original dataset. These experiments underscore our method’s ability to reduce reliance on large-scale human demonstrations while achieving robust object generalization.
Our primary contribution is a unified pipeline for integrating vision-language datasets with robot interaction data, enabling end-to-end object generalization. Through systematic evaluation, we validate the framework’s performance on complex multi-stage tasks (e.g., bin-picking) and multi-skill manipulation (e.g., rotating, pushing), highlighting its universality. Despite some of the existing works, such as RT-2 [5] and ECoT [37] giving a glimpse of how co-finetuning can achieve simple object generalization, they neither elucidate the underlying mechanism of achieving such generalization nor address the boundary of their methodologies. In contrast, our approach — though simple and straightforward — demonstrates that training VLA models with a hybrid dataset of robot interaction data and image-text data significantly enhances generalization. This level of generalization goes significantly beyond previously demonstrated end-to-end approaches. Crucially, our framework enables practical deployment: even a small set of smartphone images and brief fine-tuning suffices to adapt the model to novel objects, significantly advancing real-world robotic flexibility.
@@ 2 Related Work
Vision-language-action models for robot control. Recent research has focused on developing generalist robot policies trained on increasingly expansive robot learning datasets [26, 16, 10, 9, 21]. Vision-language-action models (VLAs) represent a promising approach for training such generalist policies [17, 38, 3, 35, 27, 25, 8, 40]. VLAs adapt vision-language models (VLMs) [41, 39, 44, 14, 32, 24, 23, 22, 1, 7], pre-trained on vast internet-scale image and text data, for robotic control. This approach offers several advantages: leveraging large vision-language model backbones, with billions of parameters, provides the necessary capacity for fitting extensive robot datasets. Furthermore, reusing weights pre-trained on internet-scale data enhances the ability of VLAs to interpret diverse language commands and generalize to novel objects and environments. However, current VLA models struggle to recognize open-world objects when these objects absent from the robot interaction data [17, 34]. This is mainly due to VLMs essentially “overwrites” its previously acquired knowledge of open-world objects with robot-specific information.
Generalization in robot learning. In the realm of robot learning, generalization, particularly object generalization, remains a core challenge and active area of research. Many works leverage techniques such as domain randomization [13], meta-learning [29, 15], retrieval-augmented generation [43], extra modality [45, 36], and data augmentation to improve a robot’s ability to recognize and interact with novel objects unseen during training. For instance, domain randomization methods [31, 13] randomize visual and physical parameters during simulation training to force the agent to learn features invariant to these irrelevant details, leading to better real-world generalization. Furthermore, meta-learning approaches [11] aim to train models that can rapidly adapt to new objects with limited data, directly addressing the object generalization problem. Finally, data augmentation methods [19, 18], enhance the diversity of the training data, exposing the model to a wider range of object appearances and orientations, thereby promoting robustness and generalization to novel objects. There is also a field of work using large language models or vision-language models to do open-vocabulary manipulation [30, 42, 20, 2], combined with motion planning and robot learning methods. However, these approaches involve separate modules that are trained independently for different components. To the best of our knowledge, this work represents the first exploration of object generalization beyond specific categories within visuomotor policy learning.
@@ 3.1 Notation and Motivation
Given a set of expert demonstrations that contain complex robot skill trajectories, we want to learn a visuomotor policy $\pi:\{\mathcal{O}_{r},\mathcal{I}_{r}\}\mapsto\mathcal{A}$ that maps the visual observations $o_{r}\in\mathcal{O}_{r}$ and the language instruction $i_{r}\in\mathcal{I}_{r}$ to actions $a\in\mathcal{A}$. The action changes accordingly when the language instruction and visual input change. The $r$ denote the data in the human demonstration data. Typically, for each language instruction it contains robot skill such as ”push” or ”pick up” and the target object, which is denoted as $\{obj_{r},skill_{r}\}\in i_{r}$. We then formally define the image-text data, where $\varphi:\{\mathcal{O}_{v},\mathcal{I}_{v}\}\mapsto\mathcal{L}_{v}$, where we input the image $o_{v}\in\mathcal{O}_{v}$ and give a language instruction $i_{v}\in\mathcal{I}_{v}$, the model is output with the corresponding answer $l_{r}\in\mathcal{L}_{v}$. The notation $v$ denotes image-text data.
In this work, we explore the generalization of objects, focusing on those that are not part of the robot interaction data but are present in image-text data.
@@ 3.2 Data Construction
image-text data construction. To explore the model’s ability to generalize to novel objects, we constructed a diverse image-text dataset. For the visual component, we collected 100 distinct objects that are not included in the robot interaction objects. Specifically, using three cameras mounted on the robot (see Figure 1), we captured 20 images per object, covering various poses and orientations to ensure diversity. For the textual component, we employ a fixed template, “Detecting the bounding box of object.”, as the question, and the corresponding bounding box as the answer. In total, our vision-language dataset comprises 2,000 image-text pairs.
Reasoning data construction. We utilize localization metadata to bridge the gap between image-text data and robot data, as previously mentioned. To establish this implicit link between image-text and action, we incorporate localization metadata into the robot data. This section details how we construct reasoning with localization for robot data.
For each task, we first identify target objects based on the language instructions. We then employ DinoX [28], a cutting-edge open-vocabulary object detector, to annotate the bounding boxes of these objects. DinoX can generate a bounding box given an object’s name. To ensure accuracy, we manually verify and correct any erroneous bounding boxes produced by DinoX. Since our workspace has two external camera views, which can result in different bounding boxes for the same object, we annotate only one (right camera in our experiments). Following Qwen2-VL [32], we use a fixed template, “<<<||||object_ref_start||||>>>{object}<<<||||object_ref_end||||>>><<<||||box_start||||>>> $(x_{1},y_{1})$,$(x_{2},y_{2})$$<$||||box_end||||$>$.”, to represent the localization reasoning. This reasoning is generated before each action and injected into the policy model through a learnable module. For a detailed explanation of this injection module’s architecture, we refer readers to DiVLA [33], the base model used in our experiments. An example of constructed image-text data is at Figure 2.
@@ 3.3 Training Strategy and Implementation Details
For our experiments, we adopt diffusion-based VLA, a widely used class of Vision-Language-Action (VLA) models exemplified by methods like $\pi_{0}$ [3] and TinyVLA [34]. We select diffusion-based VLA over auto-regressive alternatives due to its significantly faster inference speed, a critical advantage for real-time robotic applications (see FAST [27] for a detailed comparison). Specifically, we utilize DiVLA [33], a representative VLA architecture, co-train on a hybrid dataset comprising robot interaction data and the vision-language corpus. To balance task-specific adaptation and semantic generalization, we maintained a 10:1 data ratio (robot-to-image-text data) across all tasks. This ratio empirically proved sufficient for robust object generalization, aligning with prior findings on the benefits of co-training for VLA capabilities. Notably, increasing the proportion of robot data beyond this ratio led to a decline in in-domain task success rates. We hypothesize this stems from the limited capacity of the 2B-parameter DiVLA model compared to larger architectures like ECoT (7B) [37] and RT-2 (55B) [5], which can better absorb domain-specific data without overfitting.
@@ 4 Experiments
Figure 2: Example of constructed image-text data. Left: Photo taken by the robot’s camera. Right: Object captured with a smartphone.
@@ 4.1.1 Experimental Setup
Ablation study. To further understand our method’s effectiveness, we conducted an ablation study. We found that object generalization relies heavily on two key factors: first, explicitly linking vision and language to action through bounding boxes. This provides a direct connection between the visual object, its linguistic description, and the required manipulation. Second, a reasoning process for the robot data should be designed that mirrors the structure of vision-language pair data. This allows the model to leverage the rich information encoded in pre-trained vision-language models.
@@ 4.1.2 More Observations
Figure 3: Example Objects Used in Experiments. Left: Objects present in the robot training data. Right: Examples of novel objects, not present in the robot data, but included in the image-text co-training dataset (see Appendix for a comprehensive list).
@@ 4.2 Combining with More Skills
Implementation details. We train one model for each skill to ensure that the model focuses more on understanding the objects rather than multi-task learning. We use the same image-text data of “move” task. Following established protocols from our prior work, this image-text dataset trained concurrently with the demonstration data for comprehensive evaluation. Each object was tested with 3 trials. In total, 150 trials were conducted. The training setting is provided in Appendix.
Results. As shown in Table 1, our method achieved high success rates on the robot interaction objects for both rotate and push skills. Analysis of the failed “rotate” trials revealed that the primary cause is the model’s inability to grasp the target object securely. When evaluating performance on out-of-distribution (OOD) objects, we observed a decrease in task completion rates compared to in-distribution objects, as expected. However, the model still successfully completed nearly two-thirds of the trials. Notably, in most failure cases, the model did not incorrectly identify the target object but rather failed to execute the skill completely. This was particularly evident in the “rotate” trials, where successful execution hinges on a secure grasp, a challenging requirement for unseen objects. Nevertheless, these experiments strongly support the claim that ObjectVLA can transfer learned skills, beyond basic pick and place, to novel objects within the framework we have developed. The results underscore the potential of ObjectVLA for generalized robotic manipulation, capable of adapting to new objects and tasks beyond its initial training.
@@ 4.3 Instruction-Driven Bin Picking
To further evaluate ObjectVLA, we conducted experiments in a more practical scenario: end-to-end instruction-driven bin-picking. Unlike prior works (e.g., GR-2 [6] and DiVLA [33]) that execute bin-picking tasks without specific semantic instructions—typically limited to generic actions like transferring all objects from one container to another—we focus on a significantly more challenging setting [6, 33]. In our experiments, the robot is required to identify and retrieve a specific target object based on natural language instructions (e.g., ”Pick the hexagonal bolt from the bin”). This scenario elevates the complexity of conventional bin-picking tasks by integrating cross-modal understanding (vision-to-language alignment) and fine-grained object discrimination that have multiple objects in the scene. Notably, the objects are randomly placed on the panel, which is a large area. Not only does the model need to figure out the object’s position, but also needs to be aware of its pose.
Implementation details. We collected new data within this environment. For robot interaction data, we collected 600 pick-and-place trajectories using the same ”seen” objects as in previous experiments. For image-text data, we used half the number of objects from previous experiments, capturing 20 images of each. We compared our method against OpenVLA, a state-of-the-art VLA model, reporting success rates for both in-distribution and out-of-distribution objects. Evaluation consisted of three trials per object, totaling 183 trials per method. In each trial, at least two objects were randomly placed on the plate, and the model was instructed to pick and place a specific object according to the given instruction.
Results. Table 2 presents our experimental results. Bin picking, requiring object retrieval from random positions and poses, poses a significant challenge even for in-distribution objects. OpenVLA achieves a success rate of only 42.4% for in-distribution objects, significantly less than half. Surprisingly, it still completed roughly 10% of trials with out-of-distribution objects. This is likely due to some test objects sharing attributes with training objects (e.g., bread resembling a muffin, a green mug differing only in color from a brown training mug). In contrast, our method successfully completed 87 of 150 trials, including many completely novel objects, a 46.7% improvement over OpenVLA. This further emphasizes the necessity of co-training with both robot interaction and image-text data for effective object generalization.
Figure 5: Experimental results for smartphone captured objects and trained by continual learning. We test two new objects. We took pictures of these two objects via smartphone and continually trained them on a pre-trained model. Each object was evaluated across 10 trials. We report the success rate for each object.
@@ 4.4 Cheap Object Generalization via Smart-Phone Pictures and Continual Learning
The previous section demonstrated that our proposed co-training strategy enables the imitation learning method to generalize to any object by constructing corresponding image-text data for each object. This approach significantly enhances the model’s ability to handle a broader range of objects, without requiring extensive retraining. However, there are two key limitations that need to be addressed.
@@
Therefore, in this section, we test a simpler and more cost-effective approach: using a smart-phone camera to collect images from various perspectives. The model is then continuously trained on the pre-trained weights using this more accessible data collection method. As shown in Figure 5, we test with two objects: Pikachu and a brown toy cat. For each object, we capture 21 images and follow the same data construction pipeline discussed earlier. We train these objects in a bin-picking environment. Our results demonstrate that the model is able to recognize and successfully grasp the objects with a high success rate, 80% for Pikachu and 90% success rate for the toy cat. More importantly, we only need to continue training the model for 1 epoch. Because the collected data size is small, the training process can be extremely fast and can be finished up to ten minutes. This validates the effectiveness of our approach, showing that simple smartphone image collection combined with continuous learning enables open-world object manipulation in an end-to-end model. This experiment demonstrates that our method is flexible and cost-effective, making it a plug-and-play solution for existing VLA models, enabling them to generalize to virtually any object.